# TGDT: A Temporal Graph-based Digital Twin for Traffic Urban Corridors

### v1

@@ Abstract
Urban congestion at signalized intersections leads to significant delays, economic losses, and increased emissions. Existing deep learning models often lack spatial generalizability, rely on complex architectures, and struggle with real-time deployment. To address these limitations, we propose the Temporal Graph-based Digital Twin (TGDT), a scalable framework that integrates Temporal Convolutional Networks and Attentional Graph Neural Networks for dynamic, direction-aware traffic modeling and assessment at urban corridors. TGDT estimates key Measures of Effectiveness (MOEs) for traffic flow optimization at both the intersection level (e.g., queue length, waiting time) and the corridor level (e.g., traffic volume, travel time). Its modular architecture and sequential optimization scheme enable easy extension to any number of intersections and MOEs. The model outperforms state-of-the-art baselines by accurately producing high-dimensional, concurrent multi-output estimates. It also demonstrates high robustness and accuracy across diverse traffic conditions, including extreme scenarios, while relying on only a minimal set of traffic features. Fully parallelized, TGDT can simulate over a thousand scenarios within a matter of seconds, offering a cost-effective, interpretable, and real-time solution for traffic signal optimization.
@@ I Introduction
The main contributions of this paper are listed in the following:
@@
  • A sequential optimization scheme enables interdependent, hierarchical learning across TGDT’s modular architecture, improving performance.
  • TGDT offers a scalable, cost-effective solution for real-time traffic signal optimization with $O(1)$ sequential computation and GPU parallelization, simulating 1,000 scenarios in under one minute versus over 60 minutes with micro-simulators on 100 CPUs.
@@
The rest of the paper is organized as follows: Section II details the architecture of the proposed digital twin model. Section III explains the data generation methodology. Section IV evaluates the performance of our model through a series of experiments. Section V reviews related work, and finally, Section VI concludes the paper with key insights and directions for future research.
@@ II-A Sequential Optimization Technique
To support this, we define three independent Adam optimizers—each with a distinct learning rate—and use a mean squared error (MSE) loss function to guide updates for the inflow, travel time, queue length, and waiting time modules in a prioritized order. The overall loss is composed of four task-specific components, each optimized independently. Notably, the travel time loss is computed as the sum of directional losses (eastbound and westbound) across the corridor.
@@ III-A Traffic Simulation
Traffic data was generated using the SUMO (Simulation of Urban MObility) [7] micro-simulator, modeling eight east-west intersections along Florida’s SR 436 arterial road with realistic configurations and parameters. Over 50,000 hours of simulation were executed in parallel using Python’s ThreadPoolExecutor. Origin-Destination (OD) matrices were estimated using real-world ATSPM loop detector and WEJO GPS data, then used to generate routes via SUMO’s od2trips tool. The resulting dataset, Real-TMC, is based on both inferred and randomized OD matrices.
@@ IV Experimental Results
Figure 3: Visualization outputs of TGDT for a randomly selected traffic scenario. Comparison of actual (red) and predicted (green) curves for several Measures of Effectiveness (MOEs). TGDT takes ingress (inflow) traffic volumes and various parameters—such as traffic signal plans and driving behaviors—to accurately simulate bidirectional travel times throughout urban traffic corridors. It also estimates multi-directional maximum queue lengths and waiting times within a given time interval. The model first imputes intermediate inflow traffic volumes between intersections, then analyzes traffic performance across all intersections ($J_{1},...,J_{8}$) and all phase lane groups ($P_{1},...,P_{8}$) simultaneously.
@@ V Related Work
In this realm of study, Long Short-Term Memory (LSTM) networks [8] and their variants, including Bidirectional LSTM (BiLSTM) [9] and Convolutional LSTM (ConvLSTM) [6], effectively capture spatiotemporal patterns for corridor-level traffic measures. However, they lack native support for spatial dependencies and suffer from limited parallelism and vanishing gradients on long corridors. Graph Neural Networks (GNNs) [18] address spatial correlations in non-Euclidean traffic networks. Models like ST-GCN [13] and DCRNN [5] estimate travel metrics using graph-based representations, but struggle with long-term temporal modeling, input noise sensitivity, and computational demands. CNN-RNN hybrids [19] extract spatial and temporal features, yet suffer from noise sensitivity, rigid grid assumptions, and complex architectures requiring extensive tuning. Transformer-based models [12] improve temporal dynamics via attention, though they lack topological spatial awareness and demand large-scale data.
@@ VI Conclusions and Future Work
In this paper, we address the challenge of developing a deep learning framework capable of functioning as a digital twin for urban traffic corridors—providing an accurate, robust, reliable, and scalable solution for traffic performance evaluation across multiple measures and spatial-temporal scales. Our model integrates convolutional operations on multivariate time series for localized temporal encoding, applied on top of spatiotemporal representations learned through Graph Neural Networks. This enables the reconstruction of a Temporal Graph Attention Network that operates on dynamic traffic graphs with evolving edge features.
We introduce a fusion-based approach and a sequential optimization scheme that allows the model to scale effectively to any number of Measures of Effectiveness (MOEs) without compromising the accuracy of individual tasks or encountering gradient interference. The design remains cost-efficient—requiring minimal input features, and computationally optimized for parallel execution.
The model demonstrates strong resilience and robustness, as validated through a series of controlled experiments assessing its sensitivity to traffic signal cycle length, traffic volume density, and green time allocation. These experiments show the model’s effectiveness in accurately estimating bidirectional arterial travel time distributions, even under counterfactual and randomly generated traffic scenarios, using only a limited set of accessible traffic features.
For the target arterial road—Florida’s SR 436, which is 8 miles long with approximately 20-minute westbound and 15-minute eastbound travel times and includes 8 signalized intersections—the model achieves low error rates: 24 seconds for arterial travel time estimation and 100 seconds for average waiting time. It also estimates queue lengths and directional traffic volumes with a maximum error of 4 and 1.5 vehicles, respectively, at any 5-minute interval.
The framework is capable of simultaneously generating high-dimensional traffic metrics essential for corridor-level optimization. It outperforms existing solutions in both accuracy and interpretability while maintaining robustness under extreme traffic conditions and noise. This positions our model as a promising tool for real-time traffic optimization and adaptive control, contributing to the advancement of smart city infrastructure through efficient and scalable traffic management at the corridor level.
@@ VII Acknowledgments
The work was supported in part by NSF CNS 1922782. The opinions, findings, and conclusions expressed in this publication are those of the authors and not necessarily those of NSF. The authors also acknowledge the University of Florida Research Computing for providing computational resources and support that have contributed to the research results reported in this publication.

### v2

# TGDT: A Temporal Graph-based Digital Twin for Urban Traffic Corridors

@@ Abstract
Urban congestion at signalized intersections leads to significant delays, economic losses, and increased emissions. Existing deep learning models often lack spatial generalizability, rely on complex architectures, and struggle with real-time deployment. To address these limitations, we propose the Temporal Graph-based Digital Twin (TGDT), a scalable framework that integrates Temporal Convolutional Networks and Attentional Graph Neural Networks for dynamic, direction-aware traffic modeling and assessment at urban corridors. TGDT estimates key Measures of Effectiveness (MOEs) for traffic flow optimization at both the intersection level (e.g., queue length, waiting time) and the corridor level (e.g., traffic volume, travel time). Its modular architecture and sequential optimization scheme enable easy extension to any number of intersections and MOEs. The model outperforms state-of-the-art baselines by accurately producing high-dimensional, concurrent multi-output estimates. It also demonstrates high robustness and accuracy across diverse traffic conditions, including extreme scenarios, while relying on only a minimal set of traffic features. Fully parallelized, TGDT can simulate over a thousand scenarios within a matter of seconds, offering a cost-effective, interpretable, and real-time solution for urban traffic management and optimization.
@@ I Introduction
The key contributions of this paper are summarized as follows:
@@
  • A sequential optimization scheme has been employed for performance improvement, but also interdependent and hierarchical learning across TGDT’s modular architecture.
  • TGDT offers a scalable, cost-effective solution for real-time urban traffic assessment with $O(1)$ sequential computation and GPU parallelization, simulating 1,000 scenarios in under one minute versus over 60 minutes with micro-simulators on 100 CPUs.
@@
The remainder of this paper is structured as follows: Section II presents the architecture of the proposed digital twin framework. Section III describes the data generation process. Section IV provides a comprehensive evaluation of the model’s performance through experimental analysis. Section V discusses relevant prior work, and Section VI concludes the paper by summarizing key findings and outlining directions for future research.
@@ II-A Sequential Optimization Technique
To support this, we define three independent Adam optimizers with distinct learning rates and mean squared error (MSE) loss function to guide updates for the inflow, travel time, queue length, and waiting time modules in a prioritized order. The overall loss is composed of four task-specific components, each optimized independently. Notably, the travel time loss is computed as the sum of directional losses (eastbound and westbound) across the corridor.
@@ III-A Traffic Simulation
Traffic data was generated using the SUMO (Simulation of Urban MObility) [10] micro-simulator, modeling eight east-west intersections along Florida’s SR 436 arterial road with realistic configurations and parameters. Over 50,000 hours of simulation were executed in parallel using Python’s ThreadPoolExecutor. Origin-Destination (OD) matrices were estimated using real-world ATSPM loop detectors and WEJO GPS data, then used to generate routes via SUMO’s od2trips tool. The resulting dataset, Real-TMC, is based on both inferred and randomized OD matrices.
@@ IV Experimental Results
Figure 3: Visualization outputs of TGDT for a randomly selected traffic scenario. Comparison of actual (red) and predicted (green) curves for several Measures of Effectiveness (MOEs). TGDT takes ingress (inflow) traffic volumes and various parameters, such as traffic signal plans and driving behaviors, to accurately simulate bidirectional travel times throughout urban traffic corridors. It also estimates multi-directional maximum queue lengths and waiting times within a given time interval. The model first imputes intermediate inflow traffic volumes between intersections, then analyzes traffic performance across all intersections ($J_{1},...,J_{8}$) and all phase lane groups ($P_{1},...,P_{8}$) simultaneously.
@@ V Related Work
In this realm of study, Long Short-Term Memory (LSTM) networks [11] and their variants, including Bidirectional LSTM (BiLSTM) [13] and Convolutional LSTM (ConvLSTM) [9], effectively capture spatiotemporal patterns for corridor-level traffic measures. However, they lack native support for spatial dependencies and suffer from limited parallelism and vanishing gradients on long corridors.
Graph Neural Networks (GNNs) [23] address spatial correlations in non-Euclidean traffic networks. Graph convolutions are integrated with gated temporal convolutions in [23] for modeling the temporal dynamics of traffic flow in road networks, however, it falls short in adapting to dynamic graph structure in real-world urban traffic systems. Models like ST-GCN [17] and DCRNN [7] estimate travel metrics using graph-based representations, but struggle with long-term temporal modeling, input noise sensitivity, and computational demands. Graph WaveNet [18] excels at learning adaptive adjacency matrices and long-range dependencies through dilated convolutions. STGODE [2] models continuous-time dynamics using neural ordinary differential equations (ODEs) on dynamic graphs, enabling fine-grained temporal representation. TGNN [12] captures event-driven temporal transitions on dynamic graphs and is well-suited for asynchronous data. However, none of these models supports modular decomposition aligned with real-world traffic subsystems. In contrast, our TGDT employs a modular architecture inspired by digital twin principles, enabling structured multi-task outputs and allowing it to model multiple interrelated variables while maintaining structural transparency and interpretability across spatio-temporal and causal dimensions.
CNN-RNN hybrids [24] extract spatial and temporal features, i.e., [8] uses residual networks to predict citywide crowd flows by modeling complex spatial and temporal dependencies. However, they still suffer from noise sensitivity, rigid grid assumptions, and complex architectures requiring extensive tuning. Transformer-based models [16] improve temporal dynamics via attention, though they lack topological spatial awareness and demand large-scale data.
@@ VI Conclusions and Future Work
This paper tackles the challenge of designing a deep learning framework that serves as a digital twin for urban traffic corridors, offering a precise, robust, reliable, and scalable solution for evaluating traffic performance across various metrics and spatiotemporal scales. Our model integrates convolutional operations on multivariate time series for localized temporal encoding, applied on top of spatiotemporal representations learned through Graph Neural Networks. This enables the reconstruction of a Temporal Graph Attention Network that operates on dynamic traffic graphs with evolving edge features.
We introduce a fusion-based approach and a sequential optimization scheme that allows the model to scale effectively to any number of Measures of Effectiveness (MOEs) without compromising the accuracy of individual tasks or encountering gradient interference. The design remains cost-efficient, requiring minimal input features, and computationally optimized for parallel execution.
The model demonstrates strong resilience and robustness, as validated through a series of controlled experiments assessing its sensitivity to traffic signal cycle length, traffic volume density, and green time allocation. These experiments show the model’s effectiveness in accurately estimating bidirectional arterial travel time distributions, even under counterfactual and randomly generated traffic scenarios, using only a limited set of accessible traffic features. For the target arterial road (Florida’s SR 436, which is 8 miles long with approximately 20-minute westbound and 15-minute eastbound travel times and includes 8 signalized intersections), the model achieves low error rates: 24 seconds for arterial travel time estimation and 100 seconds for average waiting time. It also estimates queue lengths and directional traffic volumes with a maximum error of 4 and 1.5 vehicles, respectively, at any 5-minute interval.
We believe the introduced approach is promising in real-time traffic signal optimization and adaptive control, contributing to the advancement of smart city infrastructure through efficient and scalable traffic management at the corridor level. A promising application of this work could efficiently find optimal signal timing parameters (e.g., offset, cycle length, maximum green) across all intersections, aiming to minimize arterial travel time while constraining waiting time and queue length on minor roads, using TGDT as a surrogate deep-learning model within a Bayesian optimization [5] framework.
@@ VII Acknowledgments
The work was supported in part by NSF CNS 1922782. The opinions, findings, and conclusions expressed in this publication are those of the authors and not necessarily those of NSF.